# Neural Ordinary Differential Equations for Learning and Extrapolating System Dynamics Across Bifurcations

### v1

@@ Appendix B Hyperparameters and tuning
In addition to the tuned hyperparameters, several other parameters were held fixed throughout all experiments. These are listed in Table 3.
Table 3: Fixed hyperparameters used throughout all experiments. These were not varied during tuning.
@@
Relative ODE solver tolerance ($r_{\text{tol}}$) $1\times 10^{-5}$
@@
Batch length 20
Number of epochs 5000 (during tuning), 10000 (final experiments)
Appendix C Generating bifurcation diagrams from vector fields
To investigate the dynamical behaviour of our trained models, we used a simulation-based approach to generate the bifurcation diagrams (shown in Fig. 6). These simulations were done after training the models using the learned vector fields. For 500 different values of the bifurcation parameter $\alpha$, we ran each model until $t=2000$ from two distinct initial conditions: $(x,y)=(0,0)\nolinebreak$ and $(x,y)=(1,0.01)\nolinebreak$. From the final 500 time points of each simulation, we extracted the minimum and maximum value for the $x$-component, which serve as indicators of the system’s long-term behaviour. If the model converges to a fixed point, these values coincide; in the case of a limit cycle, they reflect the oscillation amplitude in the $x$-direction. To ensure a fair comparison, we applied the same procedure to the true system using identical initial conditions and parameter values. We restrict our bifurcation diagrams for the $x$-component, as it captures the most informative aspects of the system dynamics.
[…]
@@ Appendix C Generating bifurcation diagrams from vector fields
For Experiment 2, it can be seen that decreasing the amount of data does impact the accuracy quite a bit. There is also a clear increase in the standard deviation, indicating that the differences in performance between the seeds increases. This coincides with what we saw in Sec. 5; some random seeds were still able to reconstruct the bifurcation structure qualitatively well, but others failed. The table also shows an additional dataset, where the amount of data was increased, of which the results are discussed in Sec. D. Interestingly, doubling the amount of time series data does not lead to a big decrease in the MAE.
@@
Appendix D Additional experiment: increase data
We present an additional experiment in which the amount of data was increased to investigate the impact of data availability on model performance. Specifically, we doubled the number of sampled $\alpha$-values in the training set compared to the primary dataset. However, this increase did not lead to a substantial improvement in the model’s accuracy (Fig. 8). The additional samples failed to provide new information for the Neural ODE to further refine its learned dynamics. This suggests that the model could infer no new dynamics that were not yet represented in the primary dataset. These results highlight that, in this setting, simply increasing the amount of data does not necessarily translate to better generalization or accuracy.

### v2

@@ C.3 Hyperparameter tuning
Test case I: Lorenz
Learning rate 0.0005, 0.001, 0.005
Number of layers 1, 2, 3
Layer depth 16, 32, 64
Batch size 7, 14
Sequence length 10, 20
Test case II: Rössler
Number of layers 1, 2, 3
Layer depth 16, 32, 64
Batch size 5, 10, 25
Sequence length 10, 20, 40
Test case III: Predator-prey
@@
Table 3: Fixed hyperparameters. These were not varied during tuning.
@@
Test case I: Lorenz
ODE solver step $0.05$
Number of epochs 10000 (during tuning)
20000 (final experiments)
Test case II: Rössler
Learning rate 0.001
ODE solver step $0.1$
Number of epochs 7500 (during tuning)
15000 (final experiments)
Test case III: Predator-prey
ODE solver tolerance ($r_{\text{tol}}$) $1\times 10^{-5}$
@@
Sequence length 20
Number of epochs 5000 (during tuning)
10000 (final experiments)
Appendix D Generating bifurcation diagrams from Neural ODEs
To investigate the dynamical behaviour of our trained models, we used a simulation-based approach to generate the bifurcation diagrams that were shown in Fig.5, Fig.7 and Fig.8. These simulations were performed after training the models using the learned vector fields. We will shortly go into how the bifurcation diagrams were constructed for each of the systems.
Test case I: Lorenz
For 290 different values of the bifurcation parameter $r\in(1,30)$ we ran the model until $t=250$ from two initial conditions: $(x,y,z)=(1,1,1)\nolinebreak$ and $(x,y,z)=(-1,-1,1)\nolinebreak$. After a warm-up time, which we take to be $t=200$, we plot the $x$-value of the local minima and maxima in our trajectories. For the visualisation in Fig. 5 we used the random seed with the lowest training MAE.
Test case II: Rössler
For 300 different values of the bifurcation parameter $c\in(3,6)$ we ran the model until $t=1000$ from one initial condition: $(x,y,z)=(-4,-4,0)\nolinebreak$. After a warm-up time of $t=900$, we plot the $z$-value of the local minima and maxima in our trajectories. For the visualisation in Fig. 7 we used the random seed with the lowest training MAE.
Test case III: Predator-prey
For 500 different values of the bifurcation parameter $\alpha$, we ran each model until $t=2000$ from two distinct initial conditions: $(x,y)=(0,0)\nolinebreak$ and $(x,y)=(1,0.01)\nolinebreak$. From the final 500 time points of each simulation, we extracted the minimum and maximum value for the $x$-component, which serve as indicators of the system’s long-term behaviour. If the model converges to a fixed point, these values coincide; in the case of a limit cycle, they reflect the oscillation amplitude in the $x$-direction. To ensure a fair comparison, we applied the same procedure to the true system using identical initial conditions and parameter values. We restrict our bifurcation diagrams for the $x$-component, as it captures the most informative aspects of the system dynamics.
To determine the best performing seed we used MAE as the error metric on the simulated converged behaviour. Specifically, from the simulations described above we calculated the bifurcation error $\text{MAE}_{\mathit{bif}}$ using the minimum and maximum values for the initial condition $(x,y)=(1,0.01)$ as follows:
$\begin{aligned} &\text{MAE}_{\mathit{max}}=\frac{1}{N_{a}}\sum_{\alpha=1}^{N_{a}}\left|x_{\alpha,\mathit{max}}-\tilde{x}_{\alpha,\mathit{max}}\right|,\\ &\text{MAE}_{\mathit{min}}=\frac{1}{N_{a}}\sum_{\alpha=1}^{N_{a}}\left|x_{\alpha,\mathit{min}}-\tilde{x}_{\alpha,\mathit{min}}\right|,\\ &\text{MAE}_{\mathit{bif}}=\text{MAE}_{\mathit{max}}+\text{MAE}_{\mathit{min}},\end{aligned}$ (10)
where $N_{a}$ is the number of $\alpha$-values that was used for simulation. The $\text{MAE}_{\mathit{bif}}$ was used to select the best performing seed for each of the experiments (shown in Fig. 8).
Appendix E Extra results on test case III
Table 4 shows the mean and standard deviation of the MAE (calculated as in Eq. (10)) for each of the experiments. Note that this error metric differs from the ones used for training and validation.
Table 4: Mean $\text{MAE}_{\mathit{bif}}$ over different seeds for the different experiments. Right column is the standard deviation on the $\text{MAE}_{\mathit{bif}}$.
$\text{MAE}_{\mathit{bif}}$ Mean $\pm$ Standard deviation
@@ Appendix E Extra results on test case III
For Experiment 2, it can be seen that decreasing the amount of data does impact the accuracy quite a bit. There is also a clear increase in the standard deviation, indicating that the differences in performance between the seeds increases. This coincides with what we saw in Sec. 4; some random seeds were still able to reconstruct the bifurcation structure qualitatively well, but others failed. The table also shows an additional dataset, where the amount of data was increased, of which the results are discussed below. Interestingly, doubling the amount of time series data does not lead to a big decrease in the MAE.
@@
We also present an additional experiment in which the amount of data was increased to investigate the impact of data availability on model performance. Specifically, we doubled the number of sampled $\alpha$-values in the training set compared to the primary dataset. However, this increase did not lead to a substantial improvement in the model’s accuracy (Fig. 11). The additional samples failed to provide new information for the Neural ODE to further refine its learned dynamics. This suggests that the model could infer no new dynamics that were not yet represented in the primary dataset. These results highlight that, in this setting, simply increasing the amount of data does not necessarily translate to better generalization or accuracy.